OpenStudy (anonymous):

Should an estimate or the exact answer be calculated? Then do the appropriate calculation. Mr. Montgomery received a $20 gift card for a coffee shop. He made three purchase of $1.65 each and two more purchases of $3.95 each. How much money was then left on his card? A. An estimate is sufficient. There is about $6.00 left on the card. B. The exact answer is needed. There is $7.15 left on the card.

OpenStudy (anonymous):

MEDAL!

OpenStudy (akashdeepdeb):

$20 gift card 3 purchases of 1.65$ = 3*1.65 = 4.95$ 2 more purchases of 3.95$ = 2*3.95 = 7.9$ ^ Add these 2 to find out the total money spent. 4.95$ + 7.9$ = 12.85$ An exact answer should be calculated [Money shouldn't EVER be rounded off] So money left on Mr. Montgomery's card = 20 - 12.85$ = 7.15$ Understood? :)

OpenStudy (anonymous):

so it would be b

OpenStudy (akashdeepdeb):

yes. :)

OpenStudy (akashdeepdeb):

Which company computer do you have?

OpenStudy (anonymous):

@AkashdeepDeb :Why u asked that friend ?????:)

OpenStudy (akashdeepdeb):

Just like that! :)

OpenStudy (akashdeepdeb):

@kathy0514 Which company?

OpenStudy (anonymous):

heh !